OpenStudy (anonymous):

dont get this one help please 7x/3 + 2 = -15

OpenStudy (alfie):

You can multiply all terms by three, see what happens.

OpenStudy (alfie):

7x+6= -15*3... now solve for x.

OpenStudy (anonymous):

subtract 2 from both sides multiply by 3 then divide by 7 so x = -17(3)/7

OpenStudy (anonymous):

that isnt one of my answer :(

jhonyy9 (jhonyy9):

7x+6=-45

OpenStudy (alfie):

-51/7 is the right answer.

OpenStudy (anonymous):

lol thank u